OpenStudy (anonymous):

if k belongs to real no. and middle term of (k/2+2)^8 is 1120 . find alue of k.

OpenStudy (amistre64):

(k/2)+2 ??

OpenStudy (amistre64):

1+8+28+56+70+56+28+8+1 then for a start right?

OpenStudy (amistre64):

(k/2)^8 +8(2) (k/2)^7 +28(2)^2 (k/2)^6 +56(2)^3 (k/2)^5 +70(2)^4 (k/2)^4 +56(2)^5 (k/2)^3 +28(2)^6 (k/2)^2 +8(2)^7 (k/2) +(2)^8

OpenStudy (anonymous):

(k/2+2)=1120

OpenStudy (amistre64):

the middle terms is: 70*2^4*(k/2)^4 = 1120

OpenStudy (anonymous):

want to find value of k

OpenStudy (amistre64):

70.16.k^4/16 = 1120 70.k^4 = 1120 k^4 = 1120/70

OpenStudy (amistre64):

k = 4root(112/7) then

OpenStudy (anonymous):

oo sorry (k/2+2)^8=1120

OpenStudy (amistre64):

it said 'middle term' = 1120 ; you still want that?

OpenStudy (anonymous):

ya

OpenStudy (amistre64):

k = 4rt(16) = 2 k = +- 2; but id go with 2

OpenStudy (anonymous):

ok. thanks

OpenStudy (amistre64):

70.16.(2^4/16) = 1120 70.16 = 1120 1120 = 1120 :)